# AutoMIR: Effective Zero-Shot Medical Information Retrieval without Relevance Labels

### v1

@@ Abstract.
Medical information retrieval (MIR) is essential for retrieving relevant medical knowledge from diverse sources, including electronic health records, scientific literature, and medical databases. However, achieving effective zero-shot dense retrieval in the medical domain poses substantial challenges due to the lack of relevance-labeled data. In this paper, we introduce a novel approach called Self-Learning Hypothetical Document Embeddings (SL-HyDE) to tackle this issue. SL-HyDE leverages large language models (LLMs) as generators to generate hypothetical documents based on a given query. These generated documents encapsulate key medical context, guiding a dense retriever in identifying the most relevant documents. The self-learning framework progressively refines both pseudo-document generation and retrieval, utilizing unlabeled medical corpora without requiring any relevance-labeled data. Additionally, we present the Chinese Medical Information Retrieval Benchmark (CMIRB), a comprehensive evaluation framework grounded in real-world medical scenarios, encompassing five tasks and ten datasets. By benchmarking ten models on CMIRB, we establish a rigorous standard for evaluating medical information retrieval systems. Experimental results demonstrate that SL-HyDE significantly surpasses existing methods in retrieval accuracy while showcasing strong generalization and scalability across various LLM and retriever configurations. CMIRB data and evaluation code are publicly available at: https://github.com/CMIRB-benchmark/CMIRB
Medical Information Retrieval, Large Language Models, Zero-shot Retrieval
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Retrieval models and ranking
1. INTRODUCTION
@@
Medical information retrieval (MIR) (Luo et al., 2008; Goeuriot et al., 2016) focuses on retrieving relevant medical information from extensive data sources, such as electronic health records, scientific papers, and medical knowledge databases, based on specific medical queries. Its applications are wide-ranging, supporting doctors in clinical decision-making (Sivarajkumar et al., 2024), assisting patients in finding health information (McGowan et al., 2009), and aiding researchers in accessing relevant studies (Zheng and Yu, 2015).
In information retrieval (IR), there are two mainstream paradigms: lexical-based sparse retrieval methods, such as BM25 (Robertson et al., 2009), and embedding-based dense retrieval approaches (Xiong et al., 2020; Ding et al., 2020; Izacard et al., 2021). Dense retrievers have shown strong performance, particularly in domains where large annotated datasets are available (Karpukhin et al., 2020). These datasets are typically created by manually matching queries with relevant documents, a process that becomes impractical when dealing with collections containing millions or billions of items. As a result, many retrieval tasks face the challenge of having limited or no in-domain labeled examples.
@@
Given these challenges, several studies (Karpukhin et al., 2020; Xiong et al., 2020) have explored transfer learning for zero-shot retrieval, where dense retrievers are trained on large, high-resource datasets and evaluated on queries from different domains. MS MARCO (Bajaj et al., 2016), a dataset containing a substantial number of manually annotated query-document pairs, is commonly used for this purpose. An alternative to transfer learning is contrastive learning (Gao et al., 2021), which follows a two-stage pretraining and fine-tuning pipeline to develop general-purpose text embedding models. The pretraining stage leverages weakly supervised data gathered through large-scale web crawling, while the fine-tuning stage uses high-quality text pairs derived from data mining or manual annotation (Xiao et al., 2024; Li et al., 2023). However, the availability of such large-scale datasets or high-quality text pairs cannot always be assumed, particularly in non-English languages or specialized domains.
@@
Recently, large language models (LLMs) have demonstrated exceptional performance in zero-resource retrieval scenarios, primarily due to their extensive knowledge and robust text generation capabilities (Wang et al., 2023; Jagerman et al., 2023; Shao et al., 2023). This makes them particularly effective in situations where labeled data is scarce or unavailable. One such approach, HyDE (Hypothetical Document Embeddings)(Gao et al., 2022), employs zero-shot prompts to guide an instruction-following language model to generate hypothetical documents, effectively narrowing the semantic gap between the query and the target document. Similarly, Query2doc(Wang et al., 2023) uses few-shot prompting of LLMs to generate pseudo-documents, which are then used to expand the original query. However, applying these methods to medical information retrieval presents three critical challenges:
(1) LLMs lack the specialized medical knowledge necessary to generate highly relevant hypothetical documents. Although LLMs are trained on vast datasets drawn from a wide array of general-purpose sources, they are often insufficiently equipped with domain-specific knowledge, particularly in fields like medicine. This limitation can lead to the generation of irrelevant or misleading hypothetical documents, as LLMs may produce hallucinations—erroneous or fabricated information—due to their limited understanding of complex medical concepts.
(2) General text embedding models are inadequate for representing medical queries and documents effectively. These models are typically designed for multi-domain (e.g., medicine, law, economics) and multi-task (e.g., retrieval, question answering, classification) settings, with the aim of creating versatile retrievers capable of handling various fields, tasks, and languages. However, such generalization often fails to capture the nuanced and knowledge-intensive nature of the medical domain, where precise representation of medical terminology and context is crucial for accurate retrieval.
(3) The medical domain suffers from a scarcity of high-quality, relevance-labeled datasets. In particular, relevance-labeled medical datasets—especially in non-English languages—are extremely limited. Creating such datasets, which require the careful curation of query-document pairs, is both time-consuming and resource-intensive. The scarcity of labeled data significantly increases the cost of training and fine-tuning models to achieve high performance in medical information retrieval.
@@
In this paper, we aim to develop an effective fully zero-shot dense retrieval system that requires no relevance-labeled data for medical information retrieval. Inspired by the concept of enhancing retrieval through hypothetical documents, we design an end-to-end framework called Self-Learning Hypothetical Document Embedding (SL-HyDE). During the inference phase, the framework first employs an LLM as a generator to produce a relevant hypothetical document in response to a medical query. A retrieval model is then employed to identify the most relevant real document based on the generated hypothetical document. In the training phase, we design a self-learning mechanism that enhances the retrieval performance of SL-HyDE without the need for labeled data. Specifically, this mechanism leverages the retrieval model’s ranking capabilities to select high-relevance hypothetical documents that align with the output of the generator, effectively injecting medical knowledge into the LLM. In turn, the generator’s ability to produce high-quality hypothetical documents provides pseudo-labeled data for the retrieval model, enabling it to efficiently encode medical texts. This interactive and complementary approach generates supervisory signals that enhance both the generation and retrieval capabilities of the system. Notably, SL-HyDE begins with unlabeled medical corpora and completes the training process through a self-learning mechanism, thereby circumventing the heavy reliance on labeled data typically required for training both large language models and retrieval systems.
@@
To evaluate SL-HyDE’s performance in Chinese medical information retrieval, we identify a significant lack of suitable benchmarks. In the broader field of information retrieval, datasets like BEIR (Thakur et al., 2021) offer a collection of 18 datasets that assess embedding performance across a range of retrieval tasks. MTEB (Muennighoff et al., 2023) and C-MTEB (Xiao et al., 2024) provide extensive benchmarks for massive text embedding tasks (such as retrieval, ranking, and clustering) in English and Chinese, respectively. Despite these resources, Chinese medical datasets are significantly underrepresented. To bridge this gap, we develop a valuable Chinese Medical Information Retrieval Benchmark (CMIRB). CMIRB is constructed from real-world medical scenarios, including online consultations, medical examinations, and literature retrieval. It comprises five tasks and ten datasets, marking the first comprehensive and authentic evaluation benchmark for Chinese medical information retrieval. This benchmark is poised to accelerate advancements toward more robust and generalizable MIR systems in the future.
@@
  • We propose Self-Learning Hypothetical Document Embeddings (SL-HyDE) for zero-shot medical information retrieval, eliminating the need for relevance-labeled data.
  • We develop a comprehensive Chinese Medical Information Retrieval Benchmark (CMIRB) and evaluate the performance of various retrieval models using this benchmark.
@@
Figure 1. Training and inference pipeline of SL-HyDE.
2. RELATED WORK
2.1. Dense Retrieval
@@
In recent years, with rapid advancements in deep learning and natural language processing (NLP) technologies, researchers in the field of Information Retrieval (IR) have increasingly focused on leveraging advanced text representation methods to improve retrieval system performance. Contriever (Izacard et al., 2021) leverages unsupervised contrastive learning for dense retrieval. BGE (Xiao et al., 2024) enhances Chinese general embeddings through training on large-scale text pairs. GTE (Li et al., 2023) employs multi-stage contrastive learning for multilingual applications. mE5 (Wang et al., 2024c) uses weakly supervised pretraining and fine-tuning for multilingual embeddings. BMRETRIEVER (Xu et al., 2024) specializes in English biomedical retrieval. These works demonstrate the impact of well-structured training strategies on effective retrieval across domains. Beyond embedding-based techniques, large language models (LLMs) have demonstrated exceptional performance in zero-resource retrieval scenarios, primarily due to their extensive knowledge and robust text generation capabilities. GAR (Mao et al., 2021) enriches query semantics with generated content. HyDE (Gao et al., 2022) generates hypothetical documents for the retriever, effectively narrowing the semantic gap between the query and the target document. Query2doc (Wang et al., 2023) utilizes few-shot prompts to expand queries, boosting both sparse and dense retrieval. These methods have significantly enhanced retrieval in open-domain QA.
However, retrieval through hypothetical documents generated by LLMs often yields suboptimal results when domain-specific knowledge is insufficient. To address these challenges, we propose a self-learning framework that jointly optimizes the generator and retriever without any relevance labels, thereby enhancing retrieval performance.
2.2. Information Retrieval Benchmark
@@
To better guide the development of retrieval models, researchers have developed various datasets and benchmarks to comprehensively evaluate retrieval model performance. For instance, DuReader (He et al., 2018), a large-scale Chinese reading comprehension dataset, provides a valuable question-answer resource, significantly advancing text understanding and information retrieval research. Similarly, HotPotQA (Yang et al., 2018) introduces challenges requiring complex multi-hop reasoning, driving advancements in multi-document retrieval. BEIR(Thakur et al., 2021), a zero-shot retrieval evaluation benchmark, covers diverse retrieval tasks and offers a unified evaluation platform. MTEB (Muennighoff et al., 2023) establishes a framework for evaluating multilingual text embeddings. More recently, C-MTEB (Xiao et al., 2024) specifically addresses Chinese text embedding evaluations. Collectively, these benchmarks have propelled progress in IR technologies and provided critical resources for embedding research.
However, most benchmarks are designed for general domains, limiting their utility for specific domains such as medical retrieval. Existing medical benchmarks like CMB (Wang et al., 2024b) and CMExam (Liu et al., 2024) focus primarily on medical QA and clinical reasoning, which are not suitable for medical retrieval evaluation. To bridge this gap, we compile and adapt medical corpora specifically for IR tasks thereby improving the evaluation of medical retrievers and guiding their optimization for specialized applications.
3. METHODOLOGY
3.1. Dense Retrieval
@@
(1) $s(q,d)=<\mathcal{M}_{r}(q),\mathcal{M}_{r}(d)>$
The retrieval models then identify the top-k documents, denoted as $\overline{D}$, which have the highest similarity scores when compared to the query $q$. Earlier studies (Karpukhin et al., 2020; Ding et al., 2020) utilized dual-tower architectures to encode queries and documents separately. However, more recent research (Xiong et al., 2020; Izacard et al., 2021; Xiao et al., 2024) indicates that using the same architecture for both query and document encoding provides greater robustness in low-resource information retrieval tasks. In this work, we utilize the $\mathcal{M}_{t}$ model to serve as both the query encoder and the document encoder.
3.2. Hypothetical Document Embedding (HyDE)
@@
Large language models have achieved remarkable success in text generation across various natural language processing tasks, including question answering (Liu et al., 2021), text generation (Dathathri et al., 2019), and dialogue systems. Recently, there has been a growing interest in utilizing these models to generate relevant documents based on queries, thereby improving retrieval accuracy. Hypothetical Document Embeddings (HyDE) (Gao et al., 2022) decompose dense retrieval into two tasks: a generative task executed by an instruction-following language model and a document-document similarity task executed by a retrieval model. Initially, the query is fed to the generative model with the instruction to ”write a document that answers the question,” creating a hypothetical document. Specifically, a large language model such as ChatGPT or InstructGPT can serve as the generative model ($\mathcal{M}_{g}$). By deploying prompts, the user query is supplied, and the model generates the hypothetical documents:
(2) $d^{\prime}=\mathcal{M}_{g}(q,Prompt_{1})$
To better fuse the documents, they sample $N$ documents $[d^{\prime}_{1},...,d^{\prime}_{N}]$ from generated hypothetical documents. Subsequently, an unsupervised contrastive encoder (such as Contriever) is used to encode these documents into an embedding vector $v_{q}$.
(3) $v_{q}=\frac{1}{N+1}[\sum_{k=1}^{N}\mathcal{M}_{r}(d^{\prime}_{k})+\mathcal{M}_{r}(q)]$
The inner product is then computed between $\mathcal{M}_{r}(d)$ and the set of all document vectors:
(4) $s(q,d)=<v_{q},\mathcal{M}_{r}(d)>\forall d\in D$
This vector identifies a neighborhood in the corpus embedding space, from which similar real documents are retrieved based on vector similarity. This straightforward plug-and-play paradigm makes HyDE easy to deploy in real-world applications for solving retrieval tasks.
3.3. SL-HyDE
@@
Applying HyDE to the medical domain presents two primary challenges: (1) LLMs often lack specialized medical domain knowledge, and (2) retrievers may struggle to effectively encode medical texts due to inadequate training on medical corpora. These challenges hinder the successful implementation of HyDE technology in the medical field, making it difficult to achieve significant performance improvements in retrieval tasks. A common strategy to supplement medical domain knowledge involves fine-tuning with labeled medical data (Zhang et al., 2023; Wang et al., 2024a; Tian et al., 2023; Xu et al., 2024). However, these approaches rely on high-quality, manually constructed data to adapt general models to the medical domain. Unfortunately, obtaining such high-quality labeled data in practice is particularly challenging, making the training of a medical LLM highly costly.
@@
In this paper, we introduce a Self-Learning Hypothetical Document Embedding (SL-HyDE) mechanism designed to leverage the potential of unlabeled corpora in the medical domain. SL-HyDE consists of two main modules: a self-learning generator and a self-learning retriever. The overall framework is depicted in Figure 1.
3.4. Self-Learning Generator
A common strategy for equipping a large model with medical domain knowledge involves continued pretraining on unlabeled corpora, supervised fine-tuning on labeled data, and reinforcement learning to align the model with human preferences. However, continued pretraining typically requires substantial computational resources and extensive datasets. Consequently, supervised fine-tuning with smaller datasets has gained popularity as an efficient method for adapting large models. In this paper, we utilize the generator and retriever in SL-HyDE to create a labeled dataset based on an unsupervised corpus. The labels are entirely generated by SL-HyDE, eliminating the need for external labeled data collection.
We begin with an unlabeled corpus, specifically utilizing a medical encyclopedia dataset, Huatuo26M_encyclopedia [^1] https://huggingface.co/datasets/FreedomIntelligence/huatuo_encyclopedia_qa, as our foundational text. To construct queries, we employ a robust offline LLM, Qwen2.5222https://huggingface.co/Qwen/Qwen2.5-32B-Instruct, leveraging in-context learning (Brown, 2020; Min et al., 2022). This approach generates meaningful queries in a specific format, using prompts such as: ”Please generate a medical question based on the provided medical text. Demonstration: [DEMONSTRATION] Text: [TEXT]”.
(5) $q=LLM(d,Prompt_{2})$

### v2

@@ Abstract
Medical information retrieval (MIR) is essential for retrieving relevant medical knowledge from diverse sources, including electronic health records, scientific literature, and medical databases. However, achieving effective zero-shot dense retrieval in the medical domain poses substantial challenges due to the lack of relevance-labeled data. In this paper, we introduce a novel approach called Self-Learning Hypothetical Document Embeddings (SL-HyDE) to tackle this issue. SL-HyDE leverages large language models (LLMs) as generators to generate hypothetical documents based on a given query. These generated documents encapsulate key medical context, guiding a dense retriever in identifying the most relevant documents. The self-learning framework progressively refines both pseudo-document generation and retrieval, utilizing unlabeled medical corpora without requiring any relevance-labeled data. Additionally, we present the Chinese Medical Information Retrieval Benchmark (CMIRB), a comprehensive evaluation framework grounded in real-world medical scenarios, encompassing five tasks and ten datasets. By benchmarking ten models on CMIRB, we establish a rigorous standard for evaluating medical information retrieval systems. Experimental results demonstrate that SL-HyDE significantly surpasses HyDE in retrieval accuracy while showcasing strong generalization and scalability across various LLM and retriever configurations. Our code and data are publicly available at: https://github.com/ll0ruc/AutoMIR
AutoMIR: Effective Zero-Shot Medical Information Retrieval without Relevance Labels
Lei Li1  and Xiangxu Zhang1 and Xiao Zhou1thanks: Xiao Zhou and Zheng Liu are corresponding authors. and Zheng Liu211footnotemark: 1 1Gaoling School of Artificial Intelligence, Renmin University of China 2Beijing Academy of Artificial Intelligence {leil, xansar, xiaozhou}@ruc.edu.cn, zhengliu1026@gmail.com
1 Introduction
@@
Medical information retrieval (MIR) Luo et al. (2008); Goeuriot et al. (2016) focuses on retrieving relevant medical information from sources like electronic health records, scientific papers, and medical knowledge databases, based on specific medical queries. Its applications are wide-ranging, supporting doctors in clinical decision-making Sivarajkumar et al. (2024), assisting patients in finding health information McGowan et al. (2009), and aiding researchers in accessing relevant studies Zheng and Yu (2015).
@@
Dense retrievers Karpukhin et al. (2020); Xu et al. (2024) have shown strong performance with large labeled datasets in information retrieval (IR). Several studies Xiong et al. (2020); Li et al. (2023b); Xiao et al. (2024) have successfully employed contrastive learning to develop general-purpose text embedding models, achieving promising results in zero-resource retrieval scenarios. They leverage large-scale weakly supervised data through web crawling, or high-quality text pairs derived from data mining or manual annotation. However, the availability of such large-scale datasets cannot always be assumed, particularly in non-English languages or specialized domains.
@@
Recently, large language models (LLMs) have demonstrated exceptional performance in zero-resource retrieval scenarios Wang et al. (2023a); Shen et al. (2023); Mao et al. (2024), primarily due to their extensive knowledge and robust text generation capabilities. This makes them particularly effective in situations where labeled data is scarce or unavailable. One such approach, HyDE Gao et al. (2022), employs zero-shot prompts to guide an instruction-following language model to generate hypothetical documents, effectively narrowing the semantic gap between the query and the target document. Similarly, Query2docWang et al. (2023a) uses few-shot prompting of LLMs to generate pseudo-documents, which are then used to expand the original query. However, applying these methods to medical information retrieval presents three critical challenges: (1) LLMs lack the specialized medical knowledge necessary to generate highly relevant hypothetical documents. Although LLMs are trained on vast datasets drawn from a wide array of general-purpose sources, they are often insufficiently equipped with domain-specific knowledge, particularly in fields like medicine. (2) General text embedding models are inadequate for representing medical queries and documents effectively. These versatile retrievers are typically designed for multi-domain and multi-task settings, failing to capture the nuanced and knowledge-intensive nature of the medical domain. (3) The medical domain suffers from a scarcity of high-quality, relevance-labeled datasets. The scarcity of labeled data significantly increases the cost of training and fine-tuning these models to achieve high performance.
@@
To address these issues, we propose Self-Learning Hypothetical Document Embedding (SL-HyDE), an effective fully zero-shot dense retrieval system requiring no relevance-labeled data for medical information retrieval. During the inference phase, SL-HyDE first employs an LLM as the generator to produce a relevant hypothetical document in response to a medical query. A retrieval model is then employed to pinpoint the most relevant target document from the candidates based on the generated hypothetical document. In the training phase, we design a self-learning mechanism that enhances the retrieval performance of SL-HyDE without the need for labeled data. Specifically, this mechanism leverages the retrieval model’s ranking capabilities to select high-relevance hypothetical documents that align with the output of the generator (LLM), simultaneously injecting medical knowledge into the LLM. In turn, the generator’s ability to produce high-quality hypothetical documents provides pseudo-labeled data for the training of retrieval model, enabling it to efficiently encode medical texts. This interactive and complementary approach generates supervisory signals that enhance both the generation and retrieval capabilities of the system. Notably, SL-HyDE begins with unlabeled medical corpora and completes the training process through a self-learning mechanism, thereby circumventing the heavy reliance on labeled data typically required for training both large language models and text embedding models.
@@
To evaluate SL-HyDE’s performance in Chinese medical information retrieval, we develop a valuable Chinese Medical Information Retrieval Benchmark (CMIRB). CMIRB is constructed from real-world medical scenarios, including online consultations, medical examinations, and literature retrieval. It comprises five tasks and ten datasets, marking the first comprehensive and authentic evaluation benchmark for Chinese medical information retrieval. This benchmark is poised to accelerate advancements toward more robust and generalizable MIR systems in the future.
@@
  • We propose Self-Learning Hypothetical Document Embeddings for zero-shot medical information retrieval, eliminating the need for relevance-labeled data.
  • We are the first to develop a comprehensive Chinese Medical Information Retrieval Benchmark and evaluate the performance of various text embedding models on it.
@@
2 Related Work
2.1 Dense Retrieval
@@
Recent advancements in deep learning and natural language processing have driven improvements in information retrieval. Contriever Izacard et al. (2021) leverages unsupervised contrastive learning for dense retrieval. PEG Wu et al. (2023) and BGE Xiao et al. (2024) enhance Chinese general embeddings through training on large-scale text pairs. These works demonstrate the impact of well-structured training strategies on effective retrieval across multiple domains. Beyond embedding-based techniques, large language models have demonstrated exceptional performance in zero-resource retrieval scenarios. GAR Mao et al. (2021) enriches query semantics with generated content. HyDE Gao et al. (2022) generates hypothetical documents for the retriever, effectively narrowing the semantic gap between the query and the target document. Query2doc Wang et al. (2023a) utilizes few-shot prompts to expand queries, boosting both sparse and dense retrieval. However, retrieval through hypothetical documents generated by LLMs often yields suboptimal results when domain-specific knowledge is insufficient. To address these challenges, we propose a self-learning framework that jointly optimizes the generator and retriever without any relevance labels, thereby enhancing retrieval performance.
2.2 Information Retrieval Benchmark
@@
To better guide the development of retrieval models, researchers have developed various datasets and benchmarks. For instance, DuReader He et al. (2018), a large-scale Chinese reading comprehension dataset, significantly advances text understanding and information retrieval research. BEIRThakur et al. (2021), a zero-shot retrieval evaluation benchmark, covers diverse retrieval tasks and offers a unified evaluation platform. MTEB Muennighoff et al. (2023) establishes a framework for evaluating multilingual text embeddings. More recently, C-MTEB Xiao et al. (2024) specifically addresses Chinese text embedding evaluations. However, these benchmarks are designed for general domains, limiting their utility for specific domains such as medical retrieval. Existing medical benchmarks like CMB Wang et al. (2024b) and CMExam Liu et al. (2024) focus primarily on medical QA and clinical reasoning, which are not suitable for medical retrieval evaluation. To bridge this gap, we develop the first comprehensive and realistic evaluation benchmark based on real-world medical scenarios for Chinese medical information retrieval tasks.
3 Methodology
3.1 Preliminary
@@
$s(q,d)=<\mathcal{M}_{r}(q),\mathcal{M}_{r}(d)>.$ (1)
The retrieval models then identify the top-k documents, denoted as $D_{topk}$, which have the highest similarity scores when compared to the query $q$.
@@
Large language models have achieved remarkable success in text generation across various natural language processing tasks, including question answering Liu et al. (2021) and text generation Dathathri et al. (2019). Recently, there has been a growing interest in utilizing these models to generate relevant documents based on queries, thereby improving retrieval accuracy. Hypothetical Document Embeddings (HyDE) Gao et al. (2022) decompose dense retrieval into two tasks: a generative task executed by an instruction-following language model and a document-document similarity task executed by a retrieval model.
Figure 1: Training and inference pipeline of SL-HyDE.
3.2 Overview
@@
Applying HyDE to the medical domain presents two primary challenges: (1) LLMs often lack specialized medical domain knowledge, and (2) retrievers may struggle to effectively encode medical texts due to inadequate training on medical corpora. These challenges hinder the successful implementation of HyDE technology in the medical field, making it difficult to achieve significant performance improvements in retrieval tasks. A common strategy to supplement medical domain knowledge involves fine-tuning with labeled medical data Zhang et al. (2023); Wang et al. (2024c); Xu et al. (2024). However, these approaches rely on high-quality, manually constructed data to adapt general models to the medical domain. Unfortunately, obtaining such high-quality labeled data in practice is particularly challenging, making the training of a medical LLM highly costly.
@@
In this paper, we introduce a self-learning hypothetical document embedding mechanism designed to leverage the potential of unlabeled medical corpora. The labels are entirely generated by the generator and retriever in SL-HyDE, eliminating the need for external labeled data collection. Figure 1 presents the overall framework.
3.3 SL-HyDE Training
Self-Learning Generator. An unlabeled medical corpus, such as Huatuo26M Li et al. (2023a), serves as the foundational resource for domain-specific content. To construct queries, we employ a robust offline LLM, Qwen2.5-32B-Instruct Team (2024), leveraging in-context learning Brown (2020). With a well-designed prompt, the model effectively generates medically grounded and context-aware queries:
$q=\mathrm{LLM}(d,\mathrm{prompt}).$ (2)
To facilitate retrieval, the raw generator creates a hypothetical document that distills the relevant information from the true target document. Concretely, we provide both the query and the corresponding target document as input to the generator, along with a carefully designed prompt to guide the generation of the pseudo-document.
$d^{\prime}=\mathcal{M}_{g}(q,d,\mathrm{prompt}).$ (3)